# The discrete periodic Pitman transform: invariances, braid relations, and Burke properties

### v2

@@ A.1. Basic algebraic properties
As $\beta\to\infty$, $\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}$ converges to $\max\limits_{i}E_{i}$, uniformly as a function of the $E_{i}$ (but not as a function of $k$)
@@
First, observe that
$\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}-\max\limits_{i}E_{i}$ $=\frac{1}{\beta}\log\Bigl{(}\sum\limits_{i=0}^{k}\exp(\beta E_{i})\Bigr{)}-\frac{1}{\beta}\log\Bigl{(}\exp(\beta\max\limits_{i}E_{i})\Bigr{)}$
$=\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}.$
From here, we bound below and above.
$\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$ $\leq\frac{1}{\beta}\log\Biggl{(}\frac{k\max\limits_{i}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}=\frac{1}{\beta}\log(k).$
Similarly,
$\frac{1}{\beta}\log\Biggl{(}\frac{\sum\limits_{i=0}^{k}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}$ $\geq\frac{1}{\beta}\log\Biggl{(}\frac{\max\limits_{i}\exp(\beta E_{i})}{\max\limits_{i}\exp(\beta E_{i})}\Biggr{)}=\frac{1}{\beta}\log(1)=0.$
This concludes our proof. ∎
Lemma A.2.
The maps $\mathcal{P}$ and $\widetilde{\mathcal{P}}$ defined in (1.2) and (1.13) are involutions on $\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$.
Proof.
By [CGS24, Corollary 5.9], the map $\mathcal{W}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
$\mathcal{W}(\mathbf{X}_{1},\mathbf{X}_{2})=\bigl{(}\mathfrak{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\mathfrak{T}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr{)},$
is an involution. Therefore, combined with (A.1), for $(\mathbf{W}_{1},\mathbf{W}_{2})\in\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$, we have
$\mathcal{P}(\mathcal{P}(\mathbf{W}_{1},\mathbf{W}_{2}))$ $=(T(D(\mathbf{W}_{2},\mathbf{W}_{1}),T(\mathbf{W}_{2},\mathbf{W}_{1})),D(D(\mathbf{W}_{2},\mathbf{W}_{1}),T(\mathbf{W}_{2},\mathbf{W}_{1})))$
$=(T(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\tau_{-1}\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})),D(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\tau_{-1}\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})))$
$=(\tau_{-1}\mathfrak{T}(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\tau_{1}\tau_{-1}\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})),\mathfrak{D}(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\tau_{1}\tau_{-1}\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})))$
$=(\tau_{-1}\mathfrak{T}(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})),\mathfrak{D}(\mathfrak{D}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1}),\mathfrak{T}(\mathbf{W}_{2},\tau_{1}\mathbf{W}_{1})))$
$=(\tau_{-1}\tau_{1}\mathbf{W}_{1},\mathbf{W}_{2})$
$=(\mathbf{W}_{1},\mathbf{W}_{2}).$
Thus, $\mathcal{P}$ is an involution. Passing to zero-temperature, we get that
$\widetilde{\mathcal{P}}(\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2}))$ $=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\beta\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2}))$
$=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\mathcal{P}(\beta\mathbf{W}_{1},\beta\mathbf{W}_{2}))$
$=\lim_{\beta\to\infty}\frac{1}{\beta}(\beta\mathbf{W}_{1},\beta\mathbf{W}_{2})$
$=(\mathbf{W}_{1},\mathbf{W}_{2}),$
@@
where the penultimate equality follows from $\mathcal{P}$ being an involution and the rest follow from the uniform convergence of $\mathcal{P}$ via Lemma A.1. ∎
@@
Since $\mathcal{P}$ is an involution on $\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$, it suffices to show that $(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}\times\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$ implies $\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}_{\geq 0}^{\mathbb{Z}_{N}}$ and $\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$. Fix any $i\in\mathbb{Z}_{N}$, and take $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$. Set $Y_{i}=X_{2,i+1}-X_{1,i}$ for all $i\in\mathbb{Z}_{N}$. We observe that
@@
$\leq 0\vee\left(X_{1,i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\right)$
$\leq X_{1,i}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}.$
@@
Lemma A.4.
Let $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$. Then,
  1. (i)
    $\mathfrak{s}(T(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{1})$
  2. (ii)
    $\mathfrak{s}(D(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{2})$
  3. (iii)
    $\mathfrak{s}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{1})$
  4. (iv)
    $\mathfrak{s}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{2})$
Proof.
These all follow immediately from the definitions (1.2) and (1.13) since the sums are telescoping. ∎
Lemma A.5.
[CGS24, Lemma 5.10] Let $\mathbf{X}_{1},\mathbf{X_{2}}\in\mathbb{R}^{\mathbb{Z}_{N}}$. Then, for $i\in\mathbb{Z}_{N}$,
  1. (i)
    $D_{i}(\mathbf{X}_{1},\mathbf{X_{2}})+T_{i}(\mathbf{X}_{1},\mathbf{X_{2}})=X_{1,i}+X_{2,i}$
  2. (ii)
    $e^{-D_{i}(\mathbf{X}_{1},\mathbf{X_{2}})}+e^{-T_{i+1}(\mathbf{X}_{1},\mathbf{X_{2}})}=e^{-X_{1,i}}+e^{-X_{2,i+1}}$
  3. (iii)
Proof.
(i)-(ii) were proved in shifted form in [CGS24]. (iii) is the zero-temperature limit of (i). ∎
Lemma A.6.
[CGS24, Lemma B.3] Let $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$.
  1. (i)
    If $(\mathbf{U}_{1},\mathbf{U}_{2})=\mathcal{P}(\mathbf{X}_{1},\mathbf{X}_{2})$, then $\mathcal{P}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2})=(\tau_{1}\mathbf{U}_{1},\tau_{1}\mathbf{U}_{2})$.
  2. (ii)
    If $(\mathbf{U}_{1},\mathbf{U}_{2})=\widetilde{\mathcal{P}}(\mathbf{X}_{1},\mathbf{X}_{2})$, then $\widetilde{\mathcal{P}}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2})=(\tau_{1}\mathbf{U}_{1},\tau_{1}\mathbf{U}_{2})$.
Lemma A.7.
The determinant of the Jacobian of the transformation
$\mathcal{P}:(\mathbf{X}_{1},\mathbf{X}_{2})\to(T(\mathbf{X}_{2},\mathbf{X}_{1}),D(\mathbf{X}_{2},\mathbf{X}_{1}))$
has absolute value $1$.
Proof.
This is shown in [CGS24, Proposition 5.12]. The version of the periodic Pitman transform used there is a shifted version of the map we use, which amounts to swapping columns and rows of the Jacobian matrix, leaving the absolute value of the determinant unchanged. ∎
Lemma A.8.
For $\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{X}_{3}\in\mathbb{R}^{\mathbb{Z}_{N}}$,
$D(\mathbf{X}_{1},D(\mathbf{X}_{2},\mathbf{X}_{3}))=D\bigl{(}D(\mathbf{X}_{1},\mathbf{X}_{2}),D(T(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3}).$
Proof.
We know from [CGS24, Proposition 5.13] that
$\mathfrak{D}(\mathbf{X}_{1},\mathfrak{D}(\mathbf{X}_{2},\mathbf{X}_{3}))=\mathfrak{D}\bigl{(}\mathfrak{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\mathfrak{D}(\mathfrak{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3}\bigr{)}.$ (A.2)
By (A.1) and Lemma A.6, we have
$\tau_{1}\mathfrak{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathfrak{D}(\tau_{1}\mathbf{X}_{1},\tau_{1}\mathbf{X}_{2}).$ (A.3)
Then, we have the following:
$\quad D\bigl{(}D(\mathbf{X}_{1},\mathbf{X}_{2}),D(T(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3})\bigr{)}$
[…]
as desired. ∎
Appendix B Proofs of the zero-temperature analogues in Theorems 1.10-1.12
Because polymer models have a strong connection to last-passage percolation (LPP), our algebraic results will naturally generalize to zero-temperature equivalents, through limiting arguments. In these limits, the $(+,\times)$ algebra is swapped for the $(\max,+)$ algebra, and our maps reflect this change. As introduced in [CGS24], the ‘correct’ map to look at here is $\widetilde{D}$, the uniform limit as $\beta$ goes to infinity of $\frac{1}{\beta}D(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})$, as well as $\widetilde{T}$, which is reached through an identical limit. In this section, we use our algebraic results about the behavior of the polymer model under the periodic Pitman transform (as well as the aforementioned limiting arguments) to prove analogous facts about LPP under the zero-temperature periodic Pitman transform. The following intermediate result is needed.
Lemma B.1.
For any $k\in\mathbb{N}$ and any pair of $k$-tuples of ordered pairs $U=\bigl{(}(a_{i},b_{i})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$, $V=\bigl{(}(c_{i},d_{i})\bigr{)}_{i\in\llbracket 1,k\rrbracket}$, as $\beta\to\infty$, $\frac{1}{\beta}\log\bigl{(}Z^{\beta,\mathbf{W}}(V\mid U)\bigr{)}$ converges uniformly (as a function of $\mathbf{W}$) to $G^{\mathbf{W}}(V\mid U)$.
Proof.
This follows as a direct consequence of Lemma A.1 since there are finitely many multi-paths between any two sets of points. ∎
Note that in the special case where $U,V$ are points, this shows that the single-path polymer partition function converges to single-path LPP. Also, because the convergence to LPP is uniform on the weights, we know that the zero-temperature limit is well-behaved up to weights which depend on $\beta$. If the weights $\mathbf{W}_{\beta}$ converge to $\mathbf{\widetilde{\mathcal{W}}}$, then $\frac{1}{\beta}\log(Z^{\beta,\mathbf{W}_{\beta}}(V\mid U))$ will converge to $G^{\mathbf{\widetilde{\mathcal{W}}}}(V\mid U)$. With this in mind, we can now prove Theorem 1.10.
Proof of Theorem 1.10..
As in Theorem 1.4, we take $k_{1},k_{2},\ldots,k_{m}\in\mathbb{Z},U=\bigl{(}a_{i},b_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$, and $V=\bigl{(}(c_{i},d_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ satisfying $(U,V)\in\Psi$ and, for all $i,j\in\llbracket 1,n\rrbracket$, $a_{i}\neq k_{j}+1$ and $c_{i}\neq k_{j}$. To prove Theorem 1.10, we must show that $G^{\mathbf{W}}(V\mid U)=G^{\widetilde{\mathcal{P}}_{k_{n}}\ldots\widetilde{\mathcal{P}}_{k_{1}}\mathbf{W}}(V\mid U)$, where the operators $\mathcal{P}_{k}$ are defined in 1.9, with $T,D$ replaced by $\widetilde{T},\widetilde{D}$. Now, for $\beta>0$, let $\mathcal{P}_{k}^{\beta}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ be operators defined by their action:
$\quad\,\mathcal{P}_{k}^{\beta}(\ldots,\mathbf{W}_{k-1},\mathbf{W}_{k},\mathbf{W}_{k+1},\mathbf{W}_{k+2},\ldots)$ (B.1)
$=(\ldots,\mathbf{W}_{k-1},\frac{1}{\beta}T(\beta\mathbf{W}_{k+1},\beta\mathbf{W}_{k}),\frac{1}{\beta}D(\beta\mathbf{W}_{k+1},\beta\mathbf{W}_{k}),\mathbf{W}_{k+2},\ldots).$
By Theorem 1.4, we know that $Z^{\mathbf{W}}(V\mid U)=Z^{\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(V\mid U)$. Then, by definition of the polymer model with inverse temperature $\beta$ (1.8), for all $\beta>0$,
$Z^{\beta,\mathbf{W}}(V\mid U)=Z^{\beta\mathbf{W}}(V\mid U)=Z^{\mathcal{P}_{k_{1}}(\beta\mathbf{W})}(V\mid U)=Z^{\beta(\frac{1}{\beta}\mathcal{P}_{k_{1}}(\beta\mathbf{W}))}(V\mid U)=Z^{\beta,\mathcal{P}_{k_{1}}^{\beta}(\mathbf{W})}(V\mid U).$ (B.2)
By induction, we have $Z^{\beta,\mathbf{W}}(V\mid U)=Z^{\beta,\mathcal{P}_{k_{n}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}\mathbf{W}}(V\mid U)$.
By Lemma B, we know that $\frac{1}{\beta}\log\bigl{(}Z^{\beta,\mathbf{W}}(V\mid U)\bigr{)}$ converges to $G^{\mathbf{W}}(V\mid U)$ uniformly on the weights, so $G^{\mathbf{W}}(V\mid U)$ is preserved under the limiting operator $\lim\limits_{\beta\to\infty}\mathcal{P}_{k_{n}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}$. We claim that each $\mathcal{P}_{k}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{k}$ uniformly as a function of the weights, which implies by induction that
$G^{\mathbf{W}}(V\mid U)=G^{\lim\limits_{\beta\to\infty}\mathcal{P}_{k_{m}}^{\beta}\ldots\mathcal{P}_{k_{1}}^{\beta}\mathbf{W}}(V\mid U)=G^{\widetilde{\mathcal{P}}_{k_{m}}\ldots\widetilde{\mathcal{P}}_{k_{1}}\mathbf{W}}(V\mid U),$ (B.3)
proving the theorem. The necessary uniform convergence is shown component-wise:
$\quad\,\frac{1}{\beta}D_{i}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})-\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$
$=X_{2,i+1}+\frac{1}{\beta}\log\Biggl{(}\frac{\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i-1,j]}}}\Biggr{)}-(X_{2,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]})$
$=\Bigl{(}\frac{1}{\beta}\log\bigl{(}\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i,j]}}\bigr{)}-\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\Bigr{)}-\Bigl{(}\frac{1}{\beta}\log\bigl{(}\sum_{j\in\mathbb{Z}_{N}}e^{\beta Y_{(i-1,j]}}\bigr{)}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]}\Bigr{)}.$
By Lemma A.1, we know that $\frac{1}{\beta}\log(N)\leq\frac{1}{\beta}D_{i}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})-\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})\leq\frac{1}{\beta}\log(N)$, proving uniform convergence. An identical argument holds for $T_{i}-\widetilde{T}_{i}$. ∎
We next show the braid relations pass to zero-temperature.
Proof of Theorem 1.11..
Defining $\mathcal{P}_{k}^{\beta}$ as in (B.1), we see that $\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}=\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}$, as an immediate consequence of Theorem 1.6. As noted in the proof of Theorem 1.10, $\mathcal{P}_{k}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{k}$ uniformly. This implies the braid relation for $\widetilde{\mathcal{P}}_{k}$, since we have
To prove Theorem 1.12, we first need analogous Burke properties in the geometric and exponential cases.
Proposition B.2 (Geometric Burke property).
Take $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ to be independent sequences of independent random variables such that $X_{1,i}\sim\text{Geom}(p_{1}q_{i})$ and $X_{2,i}\sim\text{Geom}(p_{2}q_{i})$ for all $i\in\mathbb{Z}_{N}$. Then,
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})).$

### v3

@@ A.1. Additional algebraic properties
[CGS26, Proposition 5.12] The determinant of the Jacobian of the transformation
$\mathcal{P}:\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\text{ defined by }(\mathbf{X}_{1},\mathbf{X}_{2})\to\bigl(T(\mathbf{X}_{1},\mathbf{X}_{2}),D(\mathbf{X}_{1},\mathbf{X}_{2})\bigr)$
has absolute value $1$.
Lemma A.2.
[CGS26, Proposition 5.13] For $\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{X}_{3}\in\mathbb{R}^{\mathbb{Z}_{N}}$,
$D\bigl(D(\mathbf{X}_{1},\mathbf{X}_{2}),\mathbf{X}_{3}\bigr)=D\bigl(D(\mathbf{X}_{1},T(\mathbf{X}_{2},\mathbf{X}_{3})),D(\mathbf{X}_{2},\mathbf{X}_{3})).$
A.2. Proofs of the zero-temperature analogues in Theorems 1.8-1.10
Because polymer models have a strong connection to last-passage percolation (LPP), our algebraic results will naturally generalize to zero-temperature equivalents, through limiting arguments. In these limits, the $(+,\times)$ algebra is swapped for the $(\max,+)$ algebra, and our maps reflect this change. As introduced in [CGS26], the ‘correct’ map to look at here is $\widetilde{D}$, the uniform limit as $\beta$ goes to infinity of $\frac{1}{\beta}D(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})$, as well as $\widetilde{T}$, which is reached through an identical limit. In this section, we use our algebraic results about the behavior of the polymer model under the periodic Pitman transform (as well as the aforementioned limiting arguments) to prove analogous facts about LPP under the zero-temperature periodic Pitman transform. The following intermediate results are needed.
Lemma A.3.
For $y_{1},\ldots,y_{k}\in\mathbb{R}$, as $\beta\to\infty$, $\frac{1}{\beta}\log\Bigl(\sum\limits_{i=0}^{k}\exp(\beta y_{i})\Bigr)$ converges to $\max\limits_{i}y_{i}$, uniformly as a function of the $y_{i}$.
Lemma A.4.
For any $k\in\mathbb{N}$ and any pair of $k$-tuples of ordered pairs $U=\bigl((a_{i},b_{i})\bigr)_{i\in\llbracket 1,k\rrbracket}$, $V=\bigl((c_{i},d_{i})\bigr)_{i\in\llbracket 1,k\rrbracket}$, as $\beta\to\infty$, $\frac{1}{\beta}\log\bigl(Z^{\beta,\mathbf{X}}(V\mid U)\bigr)$ converges uniformly (as a function of $\mathbf{X}$) to $G^{\mathbf{X}}(V\mid U)$.
@@
This follows as a direct consequence of Lemma A.3, which is itself a standard fact, since there are only finitely many multi-paths between any two sets of points. ∎
We now prove Theorems 1.8-1.10.
Proof of Theorem 1.8..
To show that $\widetilde{\mathcal{P}}$ is an involution, we take a limit from the positive-temperature case:
$\widetilde{\mathcal{P}}(\widetilde{\mathcal{P}}(\mathbf{X}_{1},\mathbf{X}_{2}))$ $=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\beta\widetilde{\mathcal{P}}(\mathbf{X}_{1},\mathbf{X}_{2}))$
$=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\mathcal{P}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2}))$
$=\lim_{\beta\to\infty}\frac{1}{\beta}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2})$
$=(\mathbf{X}_{1},\mathbf{X}_{2}),$
@@
where the penultimate equality follows from $\mathcal{P}$ being an involution and the rest follow from the uniform convergence of $\mathcal{P}$ via Lemma A.3.
Next, for $\beta>0$, define the operators $\mathcal{P}_{k}^{\beta}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ by their action:
$\quad\,\mathcal{P}_{k}^{\beta}(\ldots,\mathbf{X}_{k-1},\mathbf{X}_{k},\mathbf{X}_{k+1},\mathbf{X}_{k+2},\ldots)$ (A.2)
$=(\ldots,\mathbf{X}_{k-1},\frac{1}{\beta}T(\beta\mathbf{X}_{k},\beta\mathbf{X}_{k+1}),\frac{1}{\beta}D(\beta\mathbf{X}_{k},\beta\mathbf{X}_{k+1}),\mathbf{X}_{k+2},\ldots).$
As an immediate consequence of Theorem 1.1(ii), we see that $\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}=\mathcal{P}_{k+1}^{\beta}\mathcal{P}_{k}^{\beta}\mathcal{P}_{k+1}^{\beta}$. By Lemma A.3, we see that $\mathcal{P}_{k}^{\beta}$ converges, as $\beta\to\infty$, to $\widetilde{\mathcal{P}}_{k}$ (uniformly as a function of the input). Then, we have
Proof of Theorem 1.9..
As in Theorem 1.4, we take $\sigma\in\mathcal{S}_{\mathbb{Z}}$ and $U=\bigl(a_{i},b_{i})_{i\in\llbracket 1,m\rrbracket}\bigr)$, and $V=\bigl((c_{i},d_{i})_{i\in\llbracket 1,m\rrbracket}\bigr)$ satisfying $(U,V)\in\mathcal{W}^{\sigma}$. To prove Theorem 1.9, we must show that $G^{\mathbf{X}}(V\mid U)=G^{\widetilde{\mathcal{P}}_{\sigma}\mathbf{X}}(V\mid U)$. By Theorem 1.4 and definition of the polymer model with parameter $\beta>0$ (1.9), we have
$Z^{\beta,\mathbf{X}}(V\mid U)=Z^{\beta\mathbf{X}}(V\mid U)=Z^{\mathcal{P}_{\sigma}(\beta\mathbf{X})}(V\mid U)=Z^{\beta(\frac{1}{\beta}\mathcal{P}_{\sigma}(\beta\mathbf{X}))}(V\mid U)=Z^{\beta,\mathcal{P}_{\sigma}^{\beta}(\mathbf{X})}(V\mid U),$ (A.3)
where in the last step, we simply define $\mathcal{P}_{\sigma}^{\beta}(\mathbf{X}):=\frac{1}{\beta}\mathcal{P}_{\sigma}(\beta\mathbf{X})$.
As mentioned in the proof of Theorem 1.8, each $\mathcal{P}_{k}^{\beta}$ converges uniformly to $\widetilde{\mathcal{P}}_{k}$ as a function of the input. By writing $\mathcal{P}_{\sigma}^{\beta}$ as a composition of these operators and using induction, we have that $\mathcal{P}_{\sigma}^{\beta}$ converges to $\widetilde{\mathcal{P}}_{\sigma}$. By Lemma A.4, we know that $\frac{1}{\beta}\log\bigl(Z^{\beta,\mathbf{X}}(V\mid U)\bigr)$ converges to $G^{\mathbf{X}}(V\mid U)$ uniformly as a function of $\mathbf{X}$. Hence, using (A.3) in the middle equality below, we have
$G^{\mathbf{X}}(V\mid U)=\lim_{\beta\to\infty}\frac{1}{\beta}\log Z^{\beta,\mathbf{X}}(V\mid U)=\lim_{\beta\to\infty}\frac{1}{\beta}\log Z^{\beta,\mathcal{P}_{\sigma}^{\beta}(\mathbf{X})}(V\mid U)=G^{\widetilde{\mathcal{P}}_{\sigma}\mathbf{X}}(V\mid U).\qed$
Before proving Theorem 1.10, we first need prove that $\widetilde{\mathcal{P}}$, restricted in its domain and range to nonnegative entries, is an involution. Then we can find analogous Burke properties in the geometric and exponential cases.
@@
Since $\widetilde{\mathcal{P}}$ is an involution on $\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ by Theorem 1.8, it suffices to show that $(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}\times\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$ implies $\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}_{\geq 0}^{\mathbb{Z}_{N}}$ and $\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$. Fix any $i\in\mathbb{Z}_{N}$, and take $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$. Set $Y_{i}=X_{1,i+1}-X_{2,i}$ for all $i\in\mathbb{Z}_{N}$. We observe that
@@
$\leq 0\vee\left(X_{1,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}\right)$
$\leq X_{1,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}.$
@@
Proposition A.6 (Geometric Burke property).
Take $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{Z}_{\geq 0}^{\mathbb{Z}_{N}}$ to be independent sequences of independent random variables such that $X_{1,i}\sim\text{Geom}(p_{1}q_{i})$ and $X_{2,i}\sim\text{Geom}(p_{2}q_{i})$ for all $i\in\mathbb{Z}_{N}$, where $p_{m}q_{i}\in(0,1)$ for $m\in\{1,2\}$ and $i\in\mathbb{Z}_{N}$. Then,
$(\mathbf{X}_{1},\mathbf{X}_{2})\overset{d}{=}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})).$